Three charged particles are held in place. Particle A has a net charge of +2 μC and is at (0, 0) m. Particle B has a net charge of +4 μC and is placed a (0.5, 0) m. Particle C has a net charge of -4 μC and is placed at (0, 0.5) m. What is the angle of direction of the net force acting on particle A?
answer: 135*
Answer #1

chuca 0.5m AL +2uc 0.5m thuc FAB = -K (8x 10-12) 0.25 FAC=k(8x10-12) 0.25 2.25 We can see fret = kx 8X10th (-1 + i) So, angle